# Caught in the act: detections of recoiling supermassive black holes from simulations

### v1

@@ 3.1 Creating mock photometric images
We begin by calculating the luminosity of the stellar particles, using stellar population synthesis results from the code Binary Population and Spectral Synthesis (BPASS, Stanway & Eldridge, 2018), and assuming a Chabrier (2003) initial mass function with mass limits $0.1\,\mathrm{M}_{\sun}$ to $100\,\mathrm{M}_{\sun}$. The age and metallicity of stellar particles is set to the median of the respective value of massive quiescent galaxies in the Large Early Galaxy Astrophysics Census (LEGA-C) presented in Bevacqua et al. (2024), namely $t_{\mathrm{age}}=3.645\,\mathrm{Gyr}$ and $Z=0.034$.
We create mock images of the galaxy merger remnant at three redshifts: a fiducal redshift of $z=0.6$, and one lower and one higher redshift: $z=0.2$ and $z=1.0$, respectively. The redshift dictates the number of pixels used in the image construction, assuming a Euclid-like spatial sampling of $\farcs 10$ 101/pixel. The observation waveband corresponds to the visible of the Euclid VIS instrument, $550\,\mathrm{nm}\mathrm{-}900\,\mathrm{nm}$ (Euclid Collaboration et al., 2024). The flux at each pixel (with the inclusion of cosmological dimming scaling as $(1+z)^{-4}$) is convolved with the Euclid transmission curves to obtain an apparent magnitude with $K$-correction. Example mock images are shown in the centre column of Figure 2 for the $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$ case, where notably the BRC is still visible.
To quantify how distinct the BRC is from the background galaxy, we calculate a non-parametric prominence measure to identify a spatially-localised intensity increase in surface brightness associated with the BRC. To do so, we use a low-pass Gaussian filter to remove high-frequency components (such as noise) from the image whilst permitting low-frequency components of the image (e.g. Davies, 1990, see Appendix A). An example of the prominence map is shown in the right panel of Figure 2. For the BRC to be detectable, we demand that the pixel hosting the BRC has a prominence $\hat{K}\geq 4.0$ (see Appendix A), translating to the pixel having such a value only once in every $10^{5}$ instances by chance alone.
@@ 3.2 Off-axis recoil and detectable fraction
To determine the amount of off-axis motion the recoiling SMBH may have and still be detected, we take from each simulation the first apocentre for each kick velocity, for all simulations where the SMBH velocity vector changes signs at least once within the simulation time of $\sim 3\,\mathrm{Gyr}$. The detectability of the BRC is determined for each snapshot for each $v_{\mathrm{kick}}$ using the prominence method in subsection 3.1. For recoil velocities $v_{\mathrm{kick}}<\sigma_{\star,0}$ the SMBH does not travel beyond the stellar core radius of the galaxy, and is undetectable. Similarly, for recoil velocities $v_{\mathrm{kick}}>1080\,\mathrm{km}\,\mathrm{s}^{-1}\simeq 0.6v_{\mathrm{esc}}$, the mass of the BRC is not enough to provide a detectable signature. Hence, we find that the minimum distance for BRC detectability in kpc, $r_{\mathrm{d}}(v_{\mathrm{kick}})$, depends linearly on $v_{\mathrm{kick}}$ as:
@@
We fit the $v_{\mathrm{kick}}$$r_{\mathrm{apo}}$ relation using a Gaussian process (GP) regressor with Stan (Stan Development Team, 2018) in the top left panel of Figure 3. To obtain a distribution of apocentres, we use in post-processing the Zlochower & Lousto (2015) model for recoil velocity that takes as input the masses, spins, and phase-space coordinates of the SMBHs prior to coalescence. We assume SMBH spins that are preferentially aligned with the SMBH binary angular momentum vector prior to merger by some angle $\phi$ (which follows a right-skewed distribution with maximum at $\phi\sim 25\fdg 0$, Lousto et al., 2010), and have a magnitude $\alpha_{\bullet}$ distribution given by
@@
which is derived from repeated SMBH binary merger simulations that follow the spin evolution of the SMBHs (Lousto et al., 2010). We then perform transformation sampling of the obtained $v_{\mathrm{kick}}$ distribution using the GP regressor to obtain a distribution of apocentre distances, from which we find that $r_{\mathrm{apo}}$ exceeds $r_{\mathrm{d}}(v_{\mathrm{kick}})$ with a detectable BRC in 33% of instances. This we take as an upper limit to the number of galaxy merger remnants with an SMBH ejected to beyond $r_{\mathrm{d}}(v_{\mathrm{kick}})$: an unrealistically optimistic case where each SMBH is ejected to an apocentre orthogonal to the LOS axis. The cumulative marginal distribution of apocentres is shown in the top right panel of Figure 3: typically, the SMBH has a maximum apocentre of $\sim 7\,\mathrm{kpc}$. The typical time to reach apocentre is $\sim 30\,\mathrm{Myr}$ following SMBH binary coalescence.
@@
For each Monte Carlo sampled value of $\tilde{v}_{\mathrm{kick}}$ in Figure 3, we weight the value of $\tilde{v}_{\mathrm{kick}}$ by its correspondingly-sampled $\cos\theta_{\mathrm{min}}$ to obtain a distribution of kick velocities weighted by observational probability, shown as blue bars in the bottom panels in Figure 3. A clear minimum is visible for $v_{\mathrm{kick}}\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$, corresponding to the extended $\theta_{\mathrm{min}}\sim 90\fdg 0$ region in the top right panel of Figure 3. This is contrasted to the steadily-decreasing $v_{\mathrm{kick}}$ probability predicted by Zlochower & Lousto (2015) for our system. Taking the cumulative sum of relative kick velocities, we find that whilst $\sim 45\%$ of recoiling SMBHs have $v_{\mathrm{kick}}\lesssim 0.6v_{\mathrm{esc}}$, only $\sim 20\%$ of all recoiling SMBHs would have a detectable BRC.
@@ 4.1 Integral field unit observations with MUSE
To confirm the presence of an SMBH within a BRC, it has been suggested (e.g. Merritt, 2013) that the velocity dispersion of the BRC would be characteristically high. However, taking stellar kinematic measurements of each individual cluster-like object in a photometric field would be observationally-expensive, and so we consider integral field unit (IFU) observations of the galaxy merger remnant to discern the imprint the recoiling SMBH leaves on the global stellar kinematics. The approach follows that done by Naab et al. (2014) and Rawlings et al. (2025) for the case where the SMBH is recoiling orthogonal to the LOS; we ignore the case of an SMBH recoiling parallel to the LOS, as the BRC would not be detected with photometric techniques.
@@
The mock IFU maps are created assuming the specifications of the Multi Unit Spectroscopic Explorer (MUSE) narrow field mode (NFM) at the Very Large Telescope, which has a field of view of $7\farcs 42$ and spatial sampling of $0\farcs 025$ per pixel (Bacon et al., 2010, see Table 1), corresponding to a physical spatial extent of $\sim 50\,\mathrm{kpc}$ and pixel width of $0.171\,\mathrm{kpc}/\mathrm{pixel}$ at our fiducial redshift of $z=0.6$. We take those stellar particles within a projected box of side length $40\,\mathrm{kpc}$, approximately the MUSE NFM field of view at our assumed redshift, and degrade the observation quality by generating 25 pseudo-particles for each stellar particle, spatially distributed as a normal distribution $\mathcal{N}(x_{i},1.372)$, where $x_{i}$ is the spatial position of stellar particle $i$ and the standard deviation is chosen as the angular resolution of the MUSE NFM instrument, both in kpc. The pseudo-particles are then binned onto a rectangular grid with $40\,\mathrm{kpc}/(0.171\,\mathrm{kpc}/\mathrm{pixel})\sim 230^{2}$ pixels, from which Voronoi bins are constructed using the method of Cappellari & Copin (2003). We require a Poisson S/N error of $N/\sqrt{N}=1000$, resulting in $10^{6}$ pseudo-particles per Voronoi bin, to which we fit a $4^{\mathrm{th}}$-order Gauss-Hermite series (Appendix B).
@@
We show the mock IFU observations for the velocity dispersion for three recoil velocities each at three different times in Figure 4. In the first row, depicting the no-recoil case, there are two distinct, symmetric ‘lobes’ of increased $\sigma$ that extend to $\sim 10\,\mathrm{kpc}$ either side of the SMBH position. The lobes persist and do not vary with time since the SMBH binary coalescence. We also create mock IFU observations in a different projection to confirm that the lobes are not an artefact of the particular viewing direction. The lobe features arise during the dynamical friction phase of the SMBH-binary merger (i.e. prior to SMBH coalescence), also causing counter-rotating regions, as mass bound to the SMBH prior to merger is lost but retains an imprint of the galaxy merger orbit (Rantala et al., 2019; Frigo et al., 2021). This feature, associated with central kinematically-decoupled regions, requires a major merger with $q\equiv M_{\star,1}/M_{\star,2}\gtrsim 1/3$ (Rantala et al., 2019).
@@
For non-zero recoil velocity, we observe systematic changes in the velocity dispersion arising from the recoiling SMBH, shown in the bottom two rows of Figure 4. The SMBH is brought to rest at apocentre, with successive apocentres decreasing due to dynamical friction with the surrounding stellar particles (Chandrasekhar, 1943).
@@ 4.2 Integral field unit observations with HARMONI
As a proof-of-concept, we also create two mock IFU observations using the specifications for the High Angular Resolution Monolithic Optical and Near-infrared Integral field spectrograph (HARMONI), a first-generation Extremely Large Telescope instrument (Thatte et al., 2010, 2016). Compared to MUSE, MICADO will boast $0\farcs 01$ spatial sampling, corresponding to a pixel width of $0.07\,\mathrm{kpc}/\mathrm{pixel}$ at the fiducial redshift $z=0.6$. We take a $1.5\,\mathrm{kpc}$ window centered on the SMBH location at an arbitrary time for the $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$ simulation (middle left panel of Figure 4) and the $v_{\mathrm{kick}}=720\,\mathrm{km}\,\mathrm{s}^{-1}$ simulation (bottom right panel in Figure 4). We again generate 25 pseudo-particles per stellar particle to create the mock IFU images, shown as the inset in the middle left and bottom right panels of Figure 4. We see a clearly elevated $\sigma$ coinciding with the BRC location, with values exceeding $240\,\mathrm{km}\,\mathrm{s}^{-1}$ in both instances. An offset peak in velocity dispersion of the order of several hundreds of kilometres per second would indicate an exciting candidate for a recoiling SMBH, making future instruments such as HARMONI ideal for such detections.
@@ 4.3 Long-slit spectroscopy
We also create mock observations of the BRC using long-slit spectroscopy (LSS). We consider three instruments: the upcoming Multi-AO Imaging Camera for Deep Observations (MICADO), another first-generation Extremely Large Telescope instrument (Sturm et al., 2024), as well as the Enhanced Resolution Imager and Spectrograph (ERIS, mounted at the VLT, e.g. Davies et al., 2023) and the recently-launched JWST NIRSPEC (e.g. Rieke et al., 2023), both currently available for observing.
@@ 4.5 Distinguishing a BRC from other objects
Figure 6: Comparison of the BRCs (coloured by $v_{\mathrm{kick}}$) to known faint objects: dwarf galaxies, ultra-compact dwarfs (UCDs), ultra-diffuse galaxies, and Milky Way globular clusters. Whilst the BRC has a similar mass and effective radius to some faint objects (left panel), it is clearly distinct from these objects in its velocity dispersion signal (right panel).
@@
In the left panel of Figure 6, increasing $v_{\mathrm{kick}}$ corresponds to a constant decrease of the effective radius of the BRC. A decrease in $M_{\mathrm{BRC}}$ with with increasing $v_{\mathrm{kick}}$ is also visible. Both trends are explained due to a higher $v_{\mathrm{kick}}$ requiring stellar particles to be more strongly bound to the BRC than a low $v_{\mathrm{kick}}$. In the $M_{\star}$-$R_{\mathrm{e}}$ plane, the BRCs mostly resemble an intermediary between dwarf systems and ultra-compact dwarfs (UCDs), that are slightly more compact than the former for a given stellar mass.
@@
In line with theoretical expectation (e.g. Merritt et al., 2009), the velocity dispersions in Figure 6 grow with increasing $v_{\mathrm{kick}}$, again owing to higher $v_{\mathrm{kick}}$ SMBHs requiring more tightly bound stellar particles to constitute the BRC.
One observed object has a velocity dispersion $\sim 100\,\mathrm{km}\,\mathrm{s}^{-1}$ and a relatively compact effective radius: this is the dwarf system M32, which has evidence for a central SMBH (van der Marel et al., 1997; McConnachie, 2012), however is unlikely to be a BRC hosting a recoiling SMBH, given the very large stellar mass of the system ($\sim 10$ times that of our simulated BRCs).

### v2

@@ 3.1 Creating mock photometric images
We begin by calculating the luminosity of the stellar particles, using stellar population synthesis results from the code Binary Population and Spectral Synthesis (BPASS, stanway2018), and assuming a chabrier2003 initial mass function with mass limits $0.1\,\mathrm{M}_{\sun}$ to $100\,\mathrm{M}_{\sun}$. The age and metallicity of stellar particles is set to the median of the respective value of massive quiescent galaxies in the Large Early Galaxy Astrophysics Census (LEGA-C) presented in bevacqua2024, namely $t_{\mathrm{age}}=3.645\,\mathrm{Gyr}$ and $Z=0.034$.
We create mock images of the galaxy merger remnant at three redshifts: a fiducal redshift of $z=0.6$, and one lower and one higher redshift: $z=0.2$ and $z=1.0$, respectively. The redshift dictates the number of pixels used in the image construction, assuming a Euclid-like spatial sampling of $0\farcs 101$/pixel. The observation waveband corresponds to the visible of the Euclid VIS instrument, $550\,\mathrm{nm}\mathrm{-}900\,\mathrm{nm}$ (cropper2024). The flux at each pixel (with the inclusion of cosmological dimming scaling as $(1+z)^{-4}$) is convolved with the Euclid transmission curves to obtain an apparent magnitude with $K$-correction. Example mock images are shown in the centre column of Figure 2 for the $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$ case, where notably the BRC is still visible.
To quantify how distinct the BRC is from the background galaxy, we calculate a non-parametric prominence measure to identify a spatially-localised intensity increase in surface brightness associated with the BRC. To do so, we use a low-pass Gaussian filter to remove high-frequency components (such as noise) from the image whilst permitting low-frequency components of the image (e.g. davies90, see Appendix A). An example of the prominence map is shown in the right panel of Figure 2. For the BRC to be detectable, we demand that the pixel hosting the BRC has a prominence $\hat{K}\geq 4.0$ (see Appendix A), translating to the pixel having such a value only once in every $10^{5}$ instances by chance alone.
@@ 3.2 Off-axis recoil and detectable fraction
To determine the amount of off-axis motion the recoiling SMBH may have and still be detected, we take from each simulation the first apocentre for each kick velocity, for all simulations where the SMBH velocity vector changes signs at least once within the simulation time of $\sim 3\,\mathrm{Gyr}$. The detectability of the BRC is determined for each snapshot for each $v_{\mathrm{kick}}$ using the prominence method in subsection 3.1. For recoil velocities $v_{\mathrm{kick}}<\sigma_{\star,0}$ the SMBH does not travel beyond the stellar core radius of the galaxy, and is undetectable. Similarly, for recoil velocities $v_{\mathrm{kick}}>1080\,\mathrm{km}\,\mathrm{s}^{-1}\simeq 0.6v_{\mathrm{esc}}$, the mass of the BRC is not enough to provide a detectable signature, in agreement with gualandris2008. Hence, we find that the minimum distance for BRC detectability in kpc, $r_{\mathrm{d}}(v_{\mathrm{kick}})$, depends linearly on $v_{\mathrm{kick}}$ as:
@@
We fit the $v_{\mathrm{kick}}$$r_{\mathrm{apo}}$ relation using a Gaussian process (GP) regressor with Stan (standevelopmentteam2018) in the top left panel of Figure 3. To obtain a distribution of apocentres, we use in post-processing the zlochower2015 model for recoil velocity that takes as input the masses, spins, and phase-space coordinates of the SMBHs prior to coalescence. We assume SMBH spins that are preferentially aligned with the SMBH binary angular momentum vector prior to merger by some angle $\phi$ (which follows a right-skewed distribution with maximum at $\phi\sim 25\fdg 0$, lousto2010), and have a magnitude $\alpha_{\bullet}$ distribution given by
@@
which is derived from repeated SMBH binary merger simulations that follow the spin evolution of the SMBHs (lousto2010). We then perform transformation sampling of the obtained $v_{\mathrm{kick}}$ distribution using the GP regressor to obtain a distribution of apocentre distances, from which we find that $r_{\mathrm{apo}}$ exceeds $r_{\mathrm{d}}(v_{\mathrm{kick}})$ with a detectable BRC in 33% of instances. This we take as an upper limit to the number of galaxy merger remnants with an SMBH ejected to beyond $r_{\mathrm{d}}(v_{\mathrm{kick}})$: an unrealistically optimistic case where each SMBH is ejected to an apocentre orthogonal to the LOS axis. The cumulative marginal distribution of apocentres is shown in the top right panel of Figure 3: typically, the SMBH has a maximum apocentre of $\sim 7\,\mathrm{kpc}$. The typical time to reach apocentre is $\sim 30\,\mathrm{Myr}$ following SMBH binary coalescence.
@@
For each Monte Carlo sampled value of $\tilde{v}_{\mathrm{kick}}$ in Figure 3, we weight the value of $\tilde{v}_{\mathrm{kick}}$ by its correspondingly-sampled $\cos\theta_{\mathrm{min}}$ to obtain a distribution of kick velocities weighted by observational probability, shown as blue bars in the bottom panels in Figure 3. A clear minimum is visible for $v_{\mathrm{kick}}\sim 600\,\mathrm{km}\,\mathrm{s}^{-1}$, corresponding to the extended $\theta_{\mathrm{min}}\sim 90\fdg 0$ region in the top right panel of Figure 3. This is contrasted to the steadily-decreasing $v_{\mathrm{kick}}$ probability predicted by zlochower2015 for our system. Taking the cumulative sum of relative kick velocities, we find that whilst $\sim 45\%$ of recoiling SMBHs have $v_{\mathrm{kick}}\lesssim 0.6v_{\mathrm{esc}}$, only $\sim 20\%$ of all recoiling SMBHs would have a detectable BRC.
We stress that the fraction of detectable BRCs discussed here are for our particular initial conditions. Whilst we do not test differing galaxy initial conditions directly, our expectation for steeper stellar density profiles would be an increase in the number of detectable BRCs at a given radius – owing to the combination of more strongly-bound stellar particles to the SMBH at the time of binary coalescence, and the more rapid decrease in the galaxy luminosity with radius. However, the deeper central potential of a galaxy with a steeper stellar density profile would likely reduce the apocentre a recoiling SMBH could achieve given some kick velocity $v_{\mathrm{kick}}$. The complex interplay of these factors makes it difficult to assess the detectable fraction of BRCs when considering different density profiles, and is an avenue for future investigation.
@@ 4.1 Integral field unit observations with MUSE
To confirm the presence of an SMBH within a BRC, it has been suggested (e.g. merritt2013) that the velocity dispersion of the BRC would be characteristically high. However, taking stellar kinematic measurements of each individual cluster-like object in a photometric field would be observationally-expensive, and so we consider integral field unit (IFU) observations of the galaxy merger remnant to discern the imprint the recoiling SMBH leaves on the global stellar kinematics. The approach follows that done by naab2014 and rawlings2025 for the case where the SMBH is recoiling orthogonal to the LOS; we ignore the case of an SMBH recoiling parallel to the LOS, as the BRC would not be detected with photometric techniques.
@@
The mock IFU maps are created assuming the specifications of the Multi Unit Spectroscopic Explorer (MUSE) narrow field mode (NFM) at the Very Large Telescope, which has a field of view of $7\farcs 42$ and spatial sampling of $0\farcs 025$ per pixel (bacon2010, see Table 1), corresponding to a physical spatial extent of $\sim 50\,\mathrm{kpc}$ and pixel width of $0.171\,\mathrm{kpc}/\mathrm{pixel}$ at our fiducial redshift of $z=0.6$. We take those stellar particles within a projected box of side length $40\,\mathrm{kpc}$, approximately the MUSE NFM field of view at our assumed redshift, and degrade the observation quality by generating 25 pseudo-particles for each stellar particle, spatially distributed as a normal distribution $\mathcal{N}(x_{i},1.372)$, where $x_{i}$ is the spatial position of stellar particle $i$ and the standard deviation is chosen as the angular resolution of the MUSE NFM instrument, both in kpc. The pseudo-particles are then binned onto a rectangular grid with $40\,\mathrm{kpc}/(0.171\,\mathrm{kpc}/\mathrm{pixel})\sim 230^{2}$ pixels, from which Voronoi bins are constructed using the method of cappellari2003. We require a Poisson S/N error of $N/\sqrt{N}=1000$, resulting in $10^{6}$ pseudo-particles per Voronoi bin, to which we fit a $4^{\mathrm{th}}$-order Gauss-Hermite series (Appendix B).
@@
We show the mock IFU observations for the velocity dispersion for three recoil velocities each at three different times in Figure 4. In the first row, depicting the no-recoil case, there are two distinct, symmetric ‘lobes’ of increased $\sigma$ that extend to $\sim 10\,\mathrm{kpc}$ either side of the SMBH position. The lobes persist and do not vary with time since the SMBH binary coalescence. We also create mock IFU observations in a different projection to confirm that the lobes are not an artefact of the particular viewing direction. The lobe features arise during the dynamical friction phase of the SMBH-binary merger (i.e. prior to SMBH coalescence), also causing counter-rotating regions, as mass bound to the SMBH prior to merger is lost but retains an imprint of the galaxy merger orbit (rantala2019; frigo2021). This feature, associated with central kinematically-decoupled regions, requires a major merger with $q\equiv M_{\star,1}/M_{\star,2}\gtrsim 1/3$ (rantala2019).
@@
For non-zero recoil velocity, we observe systematic changes in the velocity dispersion arising from the recoiling SMBH, shown in the bottom two rows of Figure 4. The SMBH is brought to rest at apocentre, with successive apocentres decreasing due to dynamical friction with the surrounding stellar particles (chandrasekhar1943).
@@ 4.2 Integral field unit observations with HARMONI
As a proof-of-concept, we also create two mock IFU observations using the specifications for the High Angular Resolution Monolithic Optical and Near-infrared Integral field spectrograph (HARMONI), a first-generation Extremely Large Telescope instrument (thatte2010; thatte2016). Compared to MUSE, MICADO will boast $0\farcs 01$ spatial sampling, corresponding to a pixel width of $0.07\,\mathrm{kpc}/\mathrm{pixel}$ at the fiducial redshift $z=0.6$. We take a $1.5\,\mathrm{kpc}$ window centered on the SMBH location at an arbitrary time for the $v_{\mathrm{kick}}=540\,\mathrm{km}\,\mathrm{s}^{-1}$ simulation (middle left panel of Figure 4) and the $v_{\mathrm{kick}}=720\,\mathrm{km}\,\mathrm{s}^{-1}$ simulation (bottom right panel in Figure 4). We again generate 25 pseudo-particles per stellar particle to create the mock IFU images, shown as the inset in the middle left and bottom right panels of Figure 4. We see a clearly elevated $\sigma$ coinciding with the BRC location, with values exceeding $240\,\mathrm{km}\,\mathrm{s}^{-1}$ in both instances. An offset peak in velocity dispersion of the order of several hundreds of kilometres per second would indicate an exciting candidate for a recoiling SMBH, making future instruments such as HARMONI ideal for such detections.
@@ 4.3 Long-slit spectroscopy
We also create mock observations of the BRC using long-slit spectroscopy (LSS). We consider three instruments: the upcoming Multi-AO Imaging Camera for Deep Observations (MICADO), another first-generation Extremely Large Telescope instrument (sturm2024), as well as the Enhanced Resolution Imager and Spectrograph (ERIS, mounted at the VLT, e.g. davies2023) and the recently-launched JWST NIRSPEC (e.g. rieke2023), both currently available for observing.
@@ 4.5 Distinguishing a BRC from other objects
Figure 6: Comparison of the BRCs (coloured by $v_{\mathrm{kick}}$) to known faint objects: dwarf galaxies, ultra-compact dwarfs (UCDs), ultra-diffuse galaxies, and Milky Way globular clusters. Whilst the BRC has a similar mass and effective radius to some faint objects (left panel), it is clearly distinct from these objects in its velocity dispersion signal (right panel). This is the case for BRCs resulting from both major (circle points) and minor (square points) mergers. For visual clarity, we only show the major merger simulations in the inset panels.
@@
In the left panel of Figure 6, increasing $v_{\mathrm{kick}}$ corresponds to a constant decrease of the effective radius of the BRC. A decrease in $M_{\mathrm{BRC}}$ with increasing $v_{\mathrm{kick}}$ is also visible – the resulting effective stellar surface mass densities of the BRCs are in the range of $\Sigma_{\mathrm{eff}}\sim 50-100\,\mathrm{M}_{\sun}\,\mathrm{pc}^{-2}$. Both trends are explained due to a higher $v_{\mathrm{kick}}$ requiring stellar particles to be more strongly-bound to the BRC than a low $v_{\mathrm{kick}}$. In the $M_{\star}$-$R_{\mathrm{e}}$ plane, the BRCs mostly resemble an intermediary between dwarf systems and ultra-compact dwarfs (UCDs), that are slightly more compact than the former for a given stellar mass.
@@
In line with theoretical expectation (e.g. merritt2009), the velocity dispersions in Figure 6 grow with increasing $v_{\mathrm{kick}}$, again owing to higher $v_{\mathrm{kick}}$ SMBHs requiring more tightly bound stellar particles to constitute the BRC.
One observed object has a velocity dispersion $\sim 100\,\mathrm{km}\,\mathrm{s}^{-1}$ and a relatively compact effective radius: this is the dwarf system M32, which has evidence for a central SMBH (vandermarel1997; mcconnachie2012), however is unlikely to be a BRC hosting a recoiling SMBH, given the very large stellar mass of the system ($\sim 10$ times that of our simulated BRCs).
@@
5 Minor mergers
Our investigation has thus far centred around the major merger of two equal-mass galaxies, and hence two equal-mass SMBHs. This introduces a bias towards larger recoil velocities that take the SMBH further from the galaxy centre than what might be expected for an unequal-mass merger. To investigate the robustness of our analysis, we consider a minor merger of mass ratio $1:5$ (where the less massive galaxy has its stellar mass reduced by a factor of 5) motivated by oser2012, who found this to be the dominant mass-weighted mass ratio for mergers involving massive galaxies. The SMBH mass and DM mass are adjusted with the stellar mass, resulting in $M_{\bullet}=5.8\times 10^{8}\,\mathrm{M}_{\sun}$ and $M_{\mathrm{DM}}=5.0\times 10^{12}\,\mathrm{M}_{\sun}$. We test recoil velocities $v_{\mathrm{kick}}=[0,420,600,900]\,\mathrm{km}\,\mathrm{s}^{-1}$. We find that the BRCs in these simulations have similar masses and velocity dispersions to the BRCs from equal-mass mergers, but are slightly more compact in effective radius (shown as the square points in Figure 6).
We find that unequal-mass mergers can produce asymmetric velocity dispersions in the mock IFU images, potentially presenting a challenge for identifying a BRC. However, any potential degeneracy between identifying a BRC and the presence of an unequal-mass merger can be broken by considering the first order asymmetric coefficient $h_{3}$ contribution to the LOSVD. In the instance of a recoiling BRC, there is no strong spatial correlation in $h_{3}$ across Voronoi cells in the mock IFU images – the $\sigma$ lobes are altered by the perturbation of a massive point-mass potential. Conversely, strong spatial correlation in $h_{3}$ is visible for unequal-mass mergers, owing to the large amount of stellar mass that alters the stellar potential as the merger remnant settles, essentially contributing a kinematically-decoupled component to the LOSVD. We show example maps in Figure 8 in Appendix C.
This does however highlight the need to consider multiple pieces of evidence for a recoiling BRC before claiming a detection, which we discuss briefly now in the context of a suggested workflow for identification.